OpenStudy (helpmeplease!!!):

If 2 x 6 = 4, 4 x 1 = 7 and 8 x 3 = 1 find the value of 5 x 5.

OpenStudy (skullpatrol):

Where did you get this question?

OpenStudy (helpmeplease!!!):

from a maths sheet

OpenStudy (skullpatrol):

Do you have an answer?

OpenStudy (anonymous):

is this maths? hahhaha

OpenStudy (skullpatrol):

Can you provide some context?

OpenStudy (skullpatrol):

Did you close this question?

OpenStudy (skullpatrol):

@HelpMePlease!!!

OpenStudy (skullpatrol):

What grade level is this question?

OpenStudy (skullpatrol):

2+6+4=12 4+1+7=12 8+3+1=12 Then 5x5=2